# RBI-ThPhys-2025-26

### v1

@@ Abstract
[…]
@@ 1 Introduction
The most recent decade has been one of many fundamental developments in cosmology, both observationally and theoretically. Notably, the advent of deep and wide galaxy surveys such as BOSS (1) and DESI (2) has provided the scientific community orders of magnitude more data on the large-scale structure (LSS) of the Universe than what was previously available. This trend will continue well with surveys such as Euclid (3), SPHEREx (4) and LSST (5) pushing to yet higher redshifts and yet more sky coverage. In order to successfully extract all cosmological and astrophysical information from large-scale structure, these developments must be accompanied by an improved theoretical understanding of the statistical properties of galaxy populations on cosmological scales.
In stark contrast to the (primary) anisotropies of the cosmic microwave background, the distribution of LSS is highly non-Gaussian due to nonlinear gravitational evolution of the initial density perturbations. Hence, the statistics of LSS are not captured by only the two-point correlation function and linear theory does not suffice except at very large separations. This has triggered a surge in the study of statistics of LSS beyond the power spectrum, the simplest of which is the bispectrum, i.e. the Fourier transform of the three-point correlation function (6). In the case of correlations between galaxy positions, the addition of such non-Gaussian statistics has been shown to yield a gain in constraining power (7; 8; 9; 10; 11).
@@
The objective of this paper and its earlier companion (25) (hereafter Paper I) is to assess the maximum possible gain in information on shape statistics by including the IA bispectrum. On the modelling side, we restrict ourselves to large scales where the bispectrum signal can be reliably computed using perturbation theory. In addition, we will assume that three-dimensional positions of galaxies are available through spectroscopic redshift measurements. In that case, the bispectrum of intrinsic alignment can be summarized in terms of multipole moments with respect to the line of sight, for which Paper I developed separable estimators. Notably, such estimators are applicable regardless of the scales being modelled (as long as a flat-sky treatment is appropriate). In Paper I, we forecasted an SNR of $\approx 30$ for $B_{DDE}$ (where $D$ denotes the galaxy density) in a DESI-like LRG sample, where shapes are measured using LSST. Direct measurements of intrinsic alignments for samples with spectroscopic redshifts can aid in isolating this effect from cosmic shear (26; 27).
@@
We use $c=1$ throughout. The cosmology used for the theory computation and simulations is Euclidean $\Lambda$CDM with $\omega_{b}=0.02225,\omega_{c}=0.1198,\Omega_{\Lambda}=0.6844,n_{s}=0.9645,\ln(10^{10}A_{s})=3.094$ which is consistent with the Planck data (29). The linear power spectrum was computed using CAMB (30).
@@ 2 Tree-level Shape Bispectrum
$g_{ij}^{\text{stoch}}(\mathbf{x})$ $=\sum_{\mathcal{O}}\epsilon_{\mathcal{O}}(\mathbf{x})\text{TF}({O}_{ij})(\mathbf{x}).$
The deterministic part, consisting of combinations of operators $\Pi_{ij}^{[n]}$ with $n=1,2$ (33), contains four operator biases $b_{1}^{\text{g}},b_{2,1}^{\text{g}},b_{2,2}^{\text{g}},b_{2,3}^{\text{g}}$. The stochastic fields possess non-trivial N-point functions for all N $>1$ and have vanishing mean (34). Some examples of the contributions arising from these fields are
@@ 2.1 Tree-level Bispectrum
×$\times B_{ijklrs}(\mathbf{k}_{1},\mathbf{k}_{2},\mathbf{k}_{3}).$
@@ 2.2 Flat Sky Projection
$\mathcal{P}^{ijkl}(\hat{\mathbf{n}})=\text{TF}(\mathcal{P}^{ik}(\hat{\mathbf{n}})\mathcal{P}^{jl}(\hat{\mathbf{n}}))$ (10)
@@
$\exp(+\mathrm{i}\phi_{\mathbf{k}})=\frac{\mathbf{k}_{x}+\mathrm{i}\mathbf{k}_{y}}{\mathbf{k}_{x}^{2}+\mathbf{k}_{y}^{2}}.$ (15)
@@
$\gamma_{E}(\mathbf{k},\hat{\mathbf{n}})$ $=\mathbf{M}^{E}_{ij}(\hat{\mathbf{k}},\hat{\mathbf{n}})^{*}g_{ij}(\mathbf{k});$ (18)
$\gamma_{B}(\mathbf{k},\hat{\mathbf{n}})$ $=\mathbf{M}^{B}_{ij}(\hat{\mathbf{k}},\hat{\mathbf{n}})^{*}g_{ij}(\mathbf{k}).$
@@
the spectra with two B-modes (that is, $B_{\delta BB}$ and $B_{EEB}$) necessarily only depend on the stochastic parameter $\alpha_{\text{gg}}^{1\delta}$; they do not depend on the $b_{2,a}^{\text{g}}$. The bispectrum consisting of three B-modes, $B_{BBB}$ vanishes identically in our model since the contribution from $\alpha_{\text{gg}}^{1\delta}$ also gets projected out. Furthermore, the deterministic piece of spectra containing any B-modes is typically strongly suppressed, and as such we do not expect these spectra to meaningfully constrain the $b_{2,a}^{\text{g}}$ bias parameters. For that reason, we will not consider the parity-odd bispectra $B_{\delta\delta B}$ and $B_{\delta EB}$ in any baseline analysis; instead we will only focus on fitting parity-even bispectra and subsequently check that the obtained results are consistent with the parity-odd sector.
@@ 2.3 Bispectrum Multipoles
The bispectrum of tensor fields admits a multipole decomposition as detailed in Paper I. There it was argued that a separable local parallel-plane Yamamoto estimator requires the use of associated Legendre polynomials to define the multipole moments (38; 39). In that work, it was also shown that differences between the use of ordinary and associated Legendre polynomials are small and that associated Legendre polynomials appear better suited for $B_{\delta\delta E}$ (at least in real space, which is the case we restrict to here). For that reason, we will only consider associated Legendre multipoles (ALMs) here. We thus also omit the superscript ‘A’ on the multipoles that was present in Paper I to distinguish between ordinary and associated Legendre multipoles, and simply refer to the ALMs as multipoles. Specifically, defining the even and odd projectors
@@
×$\times\big{(}\hat{\mathbf{n}}\cdot(\hat{\mathbf{k}}_{1}\times\hat{\mathbf{k}}_{2})\big{)},$
@@ 3 Simulations and Data Analysis
We will use a suite of 20 high-resolution dark-matter only N-body simulations from the DarkQuest Project (41). Each realization has $2048^{3}$ N-body particles with mass $M_{p}=1.02\times 10^{10}M_{\odot}/h$ and box side length $L=1\text{ Gpc}/h$. For more specifics, see e.g. (20). In this setup, we will take the halo sample in mass range $10^{12}-10^{12.5}M_{\odot}/h$ at redshift $z=0$, which we also analyzed in (28). This is especially convenient since it allows us to check whether bias parameters that are shared between the power spectrum and bispectrum of IA match. We defer performing a joint analysis to potential future work. The sample of halos in a single realization has different properties from the Luminous Red Galaxy (LRG) sample considered in Paper I. Notably, the number density is roughly $\bar{n}=3\times 10^{-3}(h/\text{Mpc})^{3}$ and $A_{\text{IA}}\approx 8$ compared to $\bar{n}=5\times 10^{-4}(h/\text{Mpc})^{3}$ and $A_{\text{IA}}\approx 4$ in Paper I, meaning that the signal-to-noise ratio (SNR) for a given total volume is much higher. This allows us to perform a more stringent test of the model under consideration.
@@ 3.1 Estimator
×$\times\hat{X}(\mathbf{q}_{1})\hat{Y}(\mathbf{q}_{2})\hat{Z}(\mathbf{q}_{3})\mathcal{P}_{XYZ}^{\ell_{1}\ell_{2}}(\{\hat{\mathbf{q}}_{i}\}),$
@@ 3.2 Covariance
Due to the limited number of realizations available to us, the sample covariance for the bispectrum is very noisy [^4] Observe that generating many independent N-body simulations to measure shapes is especially challenging; the mass resolution needs to be sufficiently high in order to reliably resolve halo shapes from their inertia tensor (45).. Thus, we are unable to probe cross-covariances between different triangle shapes, which are expected to be subdominant. However, at least for identical triangle configurations we can perform a meaningful comparison between the Gaussian approximation and the measured variance of a multipole in a given bin. We can do the same for cross-covariances between different multipoles in the same triangle bin. Formally, for two arbitrary bispectrum multipoles $B_{L},B_{L^{\prime}}$ and triangles $T,T^{\prime}$ we consider the ‘block diagonal sample variance’ estimator [^5] Moreover, when $T=T^{\prime}$ we also set all elements to zero that are expected to be zero in the Gaussian approximation. To wit, these are elements where a bispectrum without any B-modes is correlated with a bispectrum that does contain B-modes, e.g. $B^{00}_{\delta\delta E}$ and $B_{\delta BB}^{02}$. This cross-correlation is zero in the Gaussian approximation since $P_{XB}(\mathbf{k})=0$ whenever $X\neq B$.
@@
The comparison of Eq. (32) with the theoretical Gaussian covariance estimate is shown in Figure 1 for five different multipoles involving the $\delta$ and $E$ fields. For several increasing values of $k_{\text{max}}$, we consider the ratio of the sample and Gaussian covariances for all triangles with all sides below $k_{\text{max}}$ and plot the mean and sample variance of these ratios to reduce scatter. It is clear that on large scales ($k<0.10\,h$/Mpc), the Gaussian approximation works well, while for larger $k_{\text{max}}$ the Gaussian covariance (using the nonlinear matter power spectrum) tends to somewhat overpredict the error. This is likely due to our choice of using the nonlinear power spectrum rather than the linear one. Moreover, we observe at most $\sim 20\%$ offsets between the theoretical, Gaussian covariance and Eq. (32), which amounts to $\sim 10\%$ on the corresponding errors. We consider this sufficient for the purpose of this work.
Supposing that the true covariance is diagonal in triangle space (which it is in the Gaussian approximation), the estimator in Eq. (32) is unbiased. Its inverse can be computed by simply inverting, for every triangle bin $T$, the $N_{L}\times N_{L}$ block in multipole space, where $N_{L}$ is the number of multipoles. The resulting inverse is however not unbiased and needs to be ‘Hartlap-corrected’ (46). This correction needs to applied to each of the blocks, since the expectation of the inverse of the estimator is given by (47)
@@ 3.3 Discreteness and Binning Effects
×$\times(B_{\delta\delta E}^{L,\text{exact}}(T)-B_{\delta\delta E}^{L,\text{app.}}(T)).$
@@ 3.4 Performance Metrics
$b_{2,2}^{\text{g,fid}}$ $=0.129,\quad\sigma_{2,2}^{\text{fid}}=0.008;$ (40)
$b_{2,3}^{\text{g,fid}}$ $=-0.054,\quad\sigma_{2,3}^{\text{fid}}=0.017;$
@@
We will compute all the corresponding quantities define above in both the case where the Gaussian covariance is used, and the case where Eq. (32) is used, taking into account the above discussion. In order to be roughly consistent with (28), the FoB is normalized by its 68 percent confidence level. In (28), a threshold of $\Delta_{\text{crit}}=1.5$ was used to define the breakdown of the model (see also (52; 55)). In this work, given the uncertainties we face when computing the $\chi^{2}$ statistic [^13] This is due to the uncertainty in the bispectrum covariance; the large number of degrees of freedom renders the p-value for the $\chi_{\text{red}}^{2}$ statistic sensitive to even $\sim 5\%$ differences in the covariance. we opt to be somewhat more lenient and use $\Delta_{\text{crit}}=2$.
@@
As emphasized before, in order to maintain consistency with (28) we perform our baseline analysis for the volume of a single realization, i.e. $V=1\,(\text{Gpc}/h)^{3}$. When performing an MCMC analysis, uninformative uniform priors of $[-0.5,0.5]$ on the linear shape bias $b_{1}^{\text{g}}$, $[-3,3]$ on the second order bias parameters and $[-30,30]$ on the stochastic parameter $\alpha_{\text{gg}}^{1\delta}$. In the Poisson limit we would expect this parameter to be equal to
@@ 4 Results
Figure 5: Comparison of the posteriors from bispectra with at most one (green), two (grey) or three (red) shape fields at $k_{\text{max}}=0.11\,h/\text{Mpc}$. Also show is the (Gaussian) power spectrum posterior from (28), cf. Eq. (39) and (40). Observe that the grey and red posteriors are almost identical.
In our most extensive analysis, we include seven parity-even multipoles in the likelihood, to wit,
@@
Figure 7: FoB and $\chi_{\text{red}}^{2}$ for the Gaussian (blue) and diagonal sample covariance (pink) for the likelihood including bispectra with one shape field. The shaded regions indicate confidence intervals; in the top panel we plot the $1\sigma$ ($2\sigma$) confidence intervals in dark (light) grey for the FoB. In the bottom panel, the blue shades indicate the $2\sigma$ confidence intervals $[\mu_{\text{red}}-2\sigma_{\text{red}},\mu_{\text{red}}+2\sigma_{\text{red}}]$ and the pink shades $[\hat{\mu}_{\text{red}}-2\hat{\sigma}_{\text{red}},\hat{\mu}_{\text{red}}+2\hat{\sigma}_{\text{red}}]$ (see the discussion around Eq. (45)) and the dashed black lines indicate the values of $\mu_{\text{red}}=1$ and $\hat{\mu}_{\text{red}}$ as in Eq. (77).
Figure 8: Measured bispectrum signal for the four multipoles with one tensor field considered in this work. Each row contains a specific multipole, while each column contains a specific type of triangle configuration. In the left column, we plot ‘obtuse’ configurations where the sum of the two shortest sides exceeds the longest by roughly $0.02\,h/\text{Mpc}$. Note that these plots are therefore symmetric about the ‘middle bin’ for the even multipoles (first three rows); that is, the same data is plotted twice, but this is not true for $B_{\delta\delta B}^{00}$. In the leftmost and rightmost columns, we choose two specific values for the remaining ‘fixed’ side $k_{1}$ (pink and blue). The errors have been rescaled to match a volume of $V=7.5\,(\text{Gpc}/h)^{3}$.

### v2

@@ Abstract
We measure three-dimensional bispectra of halo intrinsic alignments (IA) and dark matter overdensities in real space from N-body simulations for halos of mass $10^{12}-10^{12.5}M_{\odot}/h$. We show that their multipoles with respect to the line of sight can be accurately described by a tree-level perturbation theory model on large scales ($k\lesssim 0.11\,h$/Mpc) at $z=0$. For these scales and in a simulation volume of 1 Gpc/$h$, we detect the bispectrum monopole $B_{\delta\delta E}^{00}$ at $\sim 30\sigma$ and the two quadrupoles $B_{\delta\delta E}^{11}$ and $B_{\delta\delta E}^{20}$ at $\sim 25\sigma$ and $\sim 15\sigma$, respectively. We also report similar detection significances for the lowest order multipoles of $B_{\delta EE}$ and $B_{EEE}$, although these are largely driven by stochastic contributions. We show that the first and second order EFT parameters are consistent with those obtained from fitting the IA power spectrum analysis at next-to-leading order, without requiring any priors to break degeneracies for the quadratic bias parameters. Moreover, the inclusion of higher multipole moments of $B_{\delta\delta E}$ greatly reduces the errors on second order bias parameters, by factors of 5 or more. The IA bispectrum thus provides an effective means of determining higher order shape bias parameters, thereby characterizing the scale dependence of the IA signal. We also detect parity-odd bispectra such as $B_{\delta\delta B}$ and $B_{\delta EB}$ at $\sim 10\sigma$ significance or more for k<0.15hk<0.15\,h/Mpc and they are consistent with the parity-even sector. Furthermore, we check that the Gaussian covariance approximation works reasonably well on the scales we consider here. These results lay the groundwork for using the bispectrum of IA in cosmological analyses.
@@ 1 Introduction
The most recent decade has been one of many fundamental developments in cosmology, both observationally and theoretically. Notably, the advent of deep and wide galaxy surveys such as BOSS (1) and DESI (2) has provided the scientific community orders of magnitude more data on the large-scale structure (LSS) of the Universe than what was previously available. This trend will continue well with surveys such as Euclid (3), SPHEREx (4) and LSST (5) pushing to yet higher redshifts and larger sky coverage. In order to successfully extract all cosmological and astrophysical information from large-scale structure, these developments must be accompanied by an improved theoretical understanding of the statistical properties of galaxy populations on cosmological scales.
In stark contrast to the (primary) anisotropies of the cosmic microwave background, the distribution of LSS is highly non-Gaussian due to nonlinear gravitational evolution of the initial density perturbations. Hence, the statistics of LSS are not captured by only the two-point correlation function and linear theory does not suffice except at very large separations. This has triggered a surge in the study of statistics of LSS beyond the power spectrum, the simplest of which is the bispectrum, i.e. the Fourier transform of the three-point correlation function (6). In the case of correlations between galaxy positions, the addition of such non-Gaussian statistics has been shown to yield a gain in constraining power on the cosmological model parameters (7; 8; 9; 10; 11).
@@
The objective of this paper and its earlier companion (25) (hereafter Paper I) is to assess the gain in information on shape statistics by including the IA bispectrum. On the modelling side, we restrict ourselves to large scales where the bispectrum signal can be reliably computed using perturbation theory. In addition, we will assume that three-dimensional positions of galaxies are available through spectroscopic redshift measurements. In that case, the bispectrum of intrinsic alignment can be summarized in terms of multipole moments with respect to the line of sight, for which Paper I developed separable estimators. Notably, such estimators are applicable regardless of the scales being modelled (as long as a flat-sky treatment is appropriate). In Paper I, we forecasted an SNR of $\approx 30$ for $B_{DDE}$ (where $D$ denotes the galaxy density) in a DESI-like LRG sample, where shapes are measured using LSST. Direct measurements of intrinsic alignments for samples with spectroscopic redshifts can aid in isolating this effect from cosmic shear (26; 27).
@@
[…]
@@ 2 Tree-level Shape Bispectrum
$g_{ij}^{\text{stoch}}(\mathbf{x})$ $=\sum_{\mathcal{O}}\epsilon_{\mathcal{O}}(\mathbf{x})\text{TF}(\mathcal{O}_{ij})(\mathbf{x}).$
The deterministic part, consisting of combinations of operators $\Pi_{ij}^{[n]}$ with $n=1,2$ (33), contains four operator biases $b_{1}^{\text{g}},b_{2,1}^{\text{g}},b_{2,2}^{\text{g}},b_{2,3}^{\text{g}}$. The stochastic fields possess non-trivial N-point functions for all N >1>1 and have vanishing mean (34). Some examples of the contributions arising from these fields are
@@ 2.1 Tree-level Bispectrum
$\times B_{ijklrs}(\mathbf{k}_{1},\mathbf{k}_{2},\mathbf{k}_{3}).$
@@ 2.2 Flat Sky Projection
$)\mathcal{P}^{ijkl}(\hat{\mathbf{n}})=\text{TF}(\mathcal{P}^{ik}(\hat{\mathbf{n}})\mathcal{P}^{jl}(\hat{\mathbf{n}}$)) (10)
@@
$\exp(+\mathrm{i}\phi_{\mathbf{k}})=\frac{\mathbf{k}_{x}+\mathrm{i}\mathbf{k}_{y}}{\sqrt{\mathbf{k}_{x}^{2}+\mathbf{k}_{y}^{2}}}.$ (15)
@@
[…]
@@
the spectra with two B-modes (that is, $B_{\delta BB}$ and $B_{EBB}$) necessarily only depend on the stochastic parameter $\alpha_{\text{gg}}^{1\delta}$; they do not depend on the $b_{2,a}^{\text{g}}$. The bispectrum consisting of three B-modes, $B_{BBB}$ vanishes identically in our model since the contribution from $\alpha_{\text{gg}}^{1\delta}$ also gets projected out. Furthermore, the deterministic piece of spectra containing any B-modes is typically strongly suppressed, and as such we do not expect these spectra to meaningfully constrain the $b_{2,a}^{\text{g}}$ bias parameters. For that reason, we will not consider the parity-odd bispectra $B_{\delta\delta B}$ and $B_{\delta EB}$ in any baseline analysis; instead we will only focus on fitting parity-even bispectra and subsequently check that the obtained results are broadly consistent with the parity-odd sector.
@@ 2.3 Bispectrum Multipoles
The bispectrum of tensor fields admits a multipole decomposition as detailed in Paper I. There it was argued that a separable local plane-parallel Yamamoto estimator requires the use of associated Legendre polynomials to define the multipole moments (38; 39). In that work, it was also shown that differences between the use of ordinary and associated Legendre polynomials are small and that associated Legendre polynomials appear better suited for $B_{\delta\delta E}$ (at least in real space, which is the case we restrict to here). For that reason, we will only consider associated Legendre multipoles (ALMs) here. We thus also omit the superscript ‘A’ on the multipoles that was present in Paper I to distinguish between ordinary and associated Legendre multipoles, and simply refer to the ALMs as multipoles. Specifically, defining the even and odd projectors
@@
$\times\big(\hat{\mathbf{n}}\cdot(\hat{\mathbf{k}}_{1}\times\hat{\mathbf{k}}_{2})\big),$
@@ 3 Simulations and Data Analysis
We will use a suite of 20 high-resolution dark-matter only N-body simulations from the DarkQuest Project (41). Each realization has $2048^{3}$ N-body particles with mass $M_{p}=1.02\times 10^{10}M_{\odot}/h$ and box side length $L=1\text{ Gpc}/h$. For more specifics, see e.g. (20). In this setup, we will take the halo sample in mass range $10^{12}-10^{12.5}M_{\odot}/h$ at redshift $z=0$, which we also analyzed in (28). This is especially convenient since it allows us to check whether bias parameters that are shared between the power spectrum and bispectrum of IA match. We defer performing a joint analysis to potential future work. The sample of halos in a single realization has different properties from the Luminous Red Galaxy (LRG) sample considered in Paper I. Notably, the number density is roughly $\bar{n}=3\times 10^{-3}(h/\text{Mpc})^{3}$ and $A_{\text{IA}}\approx 8$ compared to $\bar{n}=5\times 10^{-4}(h/\text{Mpc})^{3}$ and the alignment amplitude (defined in Eq. (30)) $A_{\text{IA}}\approx 4$ in Paper I, meaning that the signal-to-noise ratio (SNR) for a given total volume is much higher. This allows us to perform a more stringent test of the model under consideration.
@@ 3.1 Estimator
$\times\hat{X}(\mathbf{q}_{1})\hat{Y}(\mathbf{q}_{2})\hat{Z}(\mathbf{q}_{3})\mathcal{P}_{XYZ}^{\ell_{1}\ell_{2}}(\{\hat{\mathbf{q}}_{i}\}),$
@@ 3.2 Covariance
Due to the limited number of realizations available to us, the sample covariance for the bispectrum is very noisy [^4] Observe that generating many independent N-body simulations to measure shapes is especially challenging; the mass resolution needs to be sufficiently high in order to reliably resolve halo shapes from their inertia tensor (45).. Thus, we are unable to probe cross-covariances between different triangle shapes, which are expected to be subdominant. However, at least for identical triangle configurations we can perform a meaningful comparison between the Gaussian approximation and the measured variance of a multipole in a given bin. We can do the same for cross-covariances between different multipoles in the same triangle bin. Formally, for two arbitrary bispectrum multipoles $B_{L},B_{L^{\prime}}$ and triangles $T,T^{\prime}$ we consider the ‘block diagonal sample variance’ estimator [^5] Moreover, when $T=T^{\prime}$ we also set all elements to zero that are expected to be zero in the Gaussian approximation. These are elements where a bispectrum without any B-modes is correlated with a bispectrum that does contain B-modes, e.g. $B^{00}_{\delta\delta E}$ and $B_{\delta BB}^{02}$. This cross-correlation is zero in the Gaussian approximation since $P_{XB}(\mathbf{k})=0$ whenever $X\neq B$.
@@
The comparison of Eq. (32) with the theoretical Gaussian covariance estimate is shown in Figure 1 for five different multipoles involving the $\delta$ and $E$ fields. For several increasing values of $k_{\text{max}}$, we consider the ratio of the sample and Gaussian covariances for all triangles with all sides below $k_{\text{max}}$ and plot the mean and sample variance of these ratios to reduce scatter. It is clear that on large scales (k<0.10hk<0.10\,h/Mpc), the Gaussian approximation works well, while for larger $k_{\text{max}}$ the Gaussian covariance (using the nonlinear matter power spectrum) tends to somewhat overpredict the error. This is likely due to our choice of using the nonlinear power spectrum rather than the linear one. Moreover, we observe at most $\sim 20\%$ offsets between the theoretical, Gaussian covariance and Eq. (32), which amounts to $\sim 10\%$ on the corresponding errors. We consider this sufficient for the purpose of this work.
Supposing that the true covariance is diagonal in triangle space (as in the Gaussian approximation), the estimator in Eq. (32) is unbiased. Its inverse can be computed by simply inverting, for every triangle bin $T$, the $N_{L}\times N_{L}$ block in multipole space, where $N_{L}$ is the number of multipoles. The resulting inverse is however not unbiased and needs to be ‘Hartlap-corrected’ (46). This correction needs to applied to each of the blocks, since the expectation of the inverse of the estimator is given by (47)
@@ 3.3 Discreteness and Binning Effects
$\times(B_{\delta\delta E}^{L,\text{exact}}(T)-B_{\delta\delta E}^{L,\text{app.}}(T)).$
@@ 3.4 Performance Metrics
[…]
@@
We will compute all the corresponding quantities defined above in both the case where the Gaussian covariance is used, and the case where Eq. (32) is used, taking into account the above discussion. In order to be roughly consistent with (28), the FoB is normalized by its 68 percent confidence level. In (28), a threshold of $\Delta_{\text{crit}}=1.5$ was used to define the breakdown of the model (see also (52; 55)). In this work, given the uncertainties we face when computing the $\chi^{2}$ statistic [^13] This is due to the uncertainty in the bispectrum covariance; the large number of degrees of freedom renders the p-value for the $\chi_{\text{red}}^{2}$ statistic sensitive to even $\sim 5\%$ differences in the covariance. we opt to be somewhat more lenient and use $\Delta_{\text{crit}}=2$.
@@
As emphasized before, in order to maintain consistency with (28) we perform our baseline analysis for the volume of a single realization, i.e. $V=1\,(\text{Gpc}/h)^{3}$; that is, we fit the mean data vector and use a covariance corresponding to $V=1\,(\text{Gpc}/h)^{3}$ (see Eq. (38)). When performing an MCMC analysis, uninformative uniform priors of [0.5,0.5][-0.5,0.5] on the linear shape bias $b_{1}^{\text{g}}$, [3,3][-3,3] on the second order bias parameters and [30,30][-30,30] on the stochastic parameter $\alpha_{\text{gg}}^{1\delta}$. In the Poisson limit we would expect this parameter to be equal to
@@ 4 Results
Figure 5: Comparison of the posteriors from bispectra with at most one (green), two (grey) or three (red) shape fields at $k_{\text{max}}=0.11\,h/\text{Mpc}$. Also show is the (Gaussian) power spectrum posterior from (28), cf. Eq. (39) and (40) (where a prior on the sign of $b_{2,2}^{\text{g}}$ and $b_{2,3}^{\text{g}}$ had to be applied). Observe that the grey and red posteriors are almost identical.
In our most extensive analysis, we include seven parity-even multipoles in the likelihood,
@@
Figure 7: FoB and $\chi_{\text{red}}^{2}$ for the Gaussian (blue) and diagonal sample covariance (pink) for the likelihood including bispectra with one shape field. The shaded regions indicate confidence intervals; in the top panel we plot the $1\sigma$ ($2\sigma$) confidence intervals in dark (light) grey for the FoB. In the bottom panel, the blue shades indicate the $2\sigma$ confidence intervals $\mu_{\text{red}}-2\sigma_{\text{red}},\mu_{\text{red}}+2\sigma_{\text{red}}]$ and the pink shades $\hat{\mu}_{\text{red}}-2\hat{\sigma}_{\text{red}},\hat{\mu}_{\text{red}}+2\hat{\sigma}_{\text{red}}]$ (see the discussion around Eq. (45)) and the dashed black lines indicate the values of $\mu_{\text{red}}=1$ and $\hat{\mu}_{\text{red}}$ as in Eq. (77).
Figure 8: Measured bispectrum signal for the four multipoles with one tensor field considered in this work. We plot the product $k_{1}k_{2}k_{3}\times B$ in units of $(\text{Mpc}/h)^{3}$. Each row contains a specific multipole, while each column contains a specific type of triangle configuration. In the left column, we plot ‘obtuse’ configurations where the sum of the two shortest sides exceeds the longest by roughly $0.02\,h/\text{Mpc}$. Note that these plots are therefore symmetric about the ‘middle bin’ for the even multipoles (first three rows); that is, the same data is plotted twice, but this is not true for $B_{\delta\delta B}^{00}$. In the leftmost and rightmost columns, we choose two specific values for the remaining ‘fixed’ side $k_{1}$ (pink and blue). The errors have been rescaled to match a volume of $V=7.5\,(\text{Gpc}/h)^{3}$. The dashed lines show the theory prediction using the values from Eq. (51); we stress that these are not fits to the data.